# Quantum algorithm for partial differential equations of non-conservative systems with spatially varying parameters

### v1

@@ IV.1 Acoustic simulation
See Appendix B for the detailed derivation. The PDE (42) corresponds to Eq. (13) by setting $u(t,\bm{x})=\phi(t,\bm{x})$, $\varrho(t,\bm{x})=1/c(t,\bm{x})^{2}$, $\zeta(\bm{x})=0$, $\kappa(\bm{x})=1$, $\beta(\bm{x})=0$, $\alpha(\bm{x})=0$ and $f(\bm{x})=0$, which yield
@@
Here $\tilde{\chi}$ is the diagonal operator corresponding to the indicator function $\chi_{\Omega_{\mathrm{eval}}}(\bm{x})$ that takes $1$ if $\bm{x}\in\Omega_{\mathrm{eval}}$ and 00 otherwise, where $\Omega_{\mathrm{eval}}$ is the evaluation domain. Since the observable |$\ket{0}\!\bra{0}\otimes\tilde{c}\tilde{\chi}\tilde{c}$ is diagonal, the expectation value can be efficiently estimated by measurement outcomes of the computational basis. We would like to investigate meaningful observables in more detail and also discuss efficient estimation of their expected values in our future work.
@@ IV.2 Heat conduction simulation
Figure 4: Result of preparing MPS for coefficient oracle $O_{\mathrm{coef}}$. (a) Heat map of fidelity between the exact state $(1/\sqrt{\|\bm{c}\|_{1}})\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}\sqrt{c_{a}}\ket{a}$ and its approximation $\bm{\Phi}$ by MPS with a limited bond dimension $r_{\Phi}$ with respect to the number of integration points $2^{n_{\mathrm{a}}}$. (b) The exact amplitude $\sqrt{2^{-n_{\mathrm{frac}}}/\pi(1+k_{a}^{2})}$ and its approximation by MPS with the bond dimensions $r_{\Phi}=2$ at each integration point $k_{a}$.
@@
First, we examine the validity of the proposed MPS-based method (described in Section III.3) for implementing the LCU coefficient oracle $O_{\mathrm{coef}}$. We set the number of bond dimensions for Eq. (39) as $r_{\Psi}=10$, the tolerance parameter as $\epsilon=10^{-6}$, and the fraction part of integration points as $n_{\mathrm{frac}}=1$. Figure 4 illustrates the result of the generated MPS for approximating $O_{\mathrm{coef}}$. Figure 4(a) compares the fidelity between the exact state $(1/\sqrt{\|\bm{c}\|_{1}})\sum_{a=0}^{2^{n_{\mathrm{a}}}-1}\sqrt{c_{a}}\ket{a}$ and the MPS $\bm{\Phi}$ with the bond dimension $r_{\Phi}$; the horizontal axis is the number of integration points $2^{n_{\mathrm{a}}}$. This figure clearly shows that the bond dimension $r_{\Phi}\geq 4$ is large enough to well approximate the coefficient oracle with the fidelity greater than $0.999$. The case of the bond dimension $r_{\Phi}=2$ is also good for the relatively small number of the integration points, $2^{n_{\mathrm{anc}}}\leq 256$. For example, the fidelity was $0.98$ when $2562^{n_{\mathrm{anc}}}=$ 256 and $r_{\Phi}=2$. Figure 4(b) shows the exact amplitude $\sqrt{2^{-n_{\mathrm{frac}}}/\pi(1+k_{a}^{2})}$ and its approximation by MPS for the case of $2562^{n_{\mathrm{anc}}}=$ 256 and $r_{\Phi}=2$. These results demonstrate the validity and the effectiveness of the proposed implementation method of the coefficient oracle $O_{\mathrm{coef}}$. In the following example, we set $n_{\mathrm{anc}}=8$, $n_{\mathrm{frac}}=1$, and $r_{\Phi}=2$ because the MPS with the bond dimension of $2$ can be exactly implemented on a quantum circuit [41].
To simulate the heat equation (52), we set $n_{0}=n_{1}=4$, $\tau=1.0\times 10^{-1}$, $T=10$, $h=1$, and $\kappa=0.1$. We simulated quantum circuits by the state vector calculation. The number of qubits for encoding this problem is $n_{0}+n_{1}+n_{\mathrm{a}}=16$. As an initial state, we prepared
@@
This dynamics of the temperature describes how the initial temperature diffuses as time passes, which is a non-unitary dynamics. Note again that such direct visualization requires the full quantum state tomography and thus is impractical for a real quantum computer. Furthermore, we need to repeatedly run quantum circuits to successfully measure all zeros of the ancilla register when implementing the proposed algorithm on the actual quantum computer. In this work, we performed all the computation by statevector simulatior to verify the validity of the proposed method. We would like to conduct our future work that takes into account the sample complexity of required measurements for successfully extracting some meaningful information from the prepared time-evolved states. Also, the demonstration shown in the current work involved the two systems of PDE: one is a conservative with spatially varying parameters, and the other is non-conservative with uniform parameters, and did not tackle with non-conservative systems of PDEs with spatially varying parameters due to the computational cost although the formulation and algorithm of our proposed method include such systems. We also address this in our future work.
@@ V Conclusion
In the present study, we focused on PDEs without source terms (i.e., $\bm{b}(t)$ in Eq. (1)) for simplicity, while the original LCHS method can deal with equations including the source term, as shown in Eq. (2). Since PDEs with source terms are essential for practical applications in CAE, we plan to address this problem in our future work. Additionally, our future work will tackle the simulation of non-linear PDEs, which have already been discussed theoretically in the literature [46, 47]. We believe our current work will advance the applicability of quantum computers in the field of CAE.
@@ Acknowledgment
Appendix B Acoustic equation

### v2

@@ IV.1 Acoustic simulation
See Appendix C for the detailed derivation. The PDE (43) corresponds to Eq. (15) by setting $u(t,\bm{x})=\phi(t,\bm{x})$, $\varrho(t,\bm{x})=1/c(t,\bm{x})^{2}$, $\zeta(\bm{x})=0$, $\kappa(\bm{x})=1$, $\beta(\bm{x})=0$, $\alpha(\bm{x})=0$ and $f(\bm{x})=0$, which yield
@@
Here $\tilde{\chi}$ is the diagonal operator corresponding to the indicator function $\chi_{\Omega_{\mathrm{eval}}}(\bm{x})$ that takes $1$ if $\bm{x}\in\Omega_{\mathrm{eval}}$ and 00 otherwise, where $\Omega_{\mathrm{eval}}$ is the evaluation domain. Since the observable |$\ket{0}\!\bra{0}\otimes\tilde{c}\tilde{\chi}\tilde{c}$ is diagonal, the expectation value can be efficiently estimated by measurement outcomes of the computational basis. In this way, if one is interested in a local quantity at a certain subdomain, one can evaluate such a quantity efficiently using the diagonal operator corresponding to the indicator function. The power spectra [18] is also a possible observable for an acoustic system. We would like to investigate meaningful observables in more detail and also discuss efficient estimation of their expected values in our future work.
@@ IV.2 Heat conduction simulation
Figure 6: Result of preparing MPS for coefficient oracle $O_{\mathrm{coef}}$. (a) Heat map of fidelity between the exact state $(1/\sqrt{\|\bm{c}\|_{1}})\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}\sqrt{c_{a}}\ket{a}$ and its approximation $\bm{\Phi}$ by MPS with a limited bond dimension $r_{\Phi}$ with respect to the number of integration points $2^{n_{\mathrm{anc}}}$. (b) The exact amplitude $\sqrt{2^{-n_{\mathrm{frac}}}/\pi(1+k_{a}^{2})}$ and its approximation by MPS with the bond dimensions $r_{\Phi}=2$ at each integration point $k_{a}$.
@@
First, we examine the validity of the proposed MPS-based method (described in Section III.3) for implementing the LCU coefficient oracle $O_{\mathrm{coef}}$. We set the number of bond dimensions for Eq. (41) as $r_{\Psi}=10$, the tolerance parameter as $\epsilon=10^{-6}$, and the fraction part of integration points as $n_{\mathrm{frac}}=1$. Figure 6 illustrates the result of the generated MPS for approximating $O_{\mathrm{coef}}$. Figure 6(a) compares the fidelity between the exact state $(1/\sqrt{\|\bm{c}\|_{1}})\sum_{a=0}^{2^{n_{\mathrm{anc}}}-1}\sqrt{c_{a}}\ket{a}$ and the MPS $\bm{\Phi}$ with the bond dimension $r_{\Phi}$; the horizontal axis is the number of integration points $2^{n_{\mathrm{anc}}}$. This figure clearly shows that the bond dimension $r_{\Phi}\geq 4$ is large enough to well approximate the coefficient oracle with the fidelity greater than $0.999$. The case of the bond dimension $r_{\Phi}=2$ is also good for the relatively small number of the integration points, $2^{n_{\mathrm{anc}}}\leq 256$. For example, the fidelity was $0.98$ when $2562^{n_{\mathrm{anc}}}=$ 256 and $r_{\Phi}=2$. Figure 6(b) shows the exact amplitude $\sqrt{2^{-n_{\mathrm{frac}}}/\pi(1+k_{a}^{2})}$ and its approximation by MPS for the case of $2562^{n_{\mathrm{anc}}}=$ 256 and $r_{\Phi}=2$. These results demonstrate the validity and the effectiveness of the proposed implementation method of the coefficient oracle $O_{\mathrm{coef}}$. Since the coefficient oracle is not problem specific but is common for the implementation of LCHS, we confirm that our proposed method will work well for various applications based on LCHS. In the following example, we set $n_{\mathrm{anc}}=8$, $n_{\mathrm{frac}}=1$, and $r_{\Phi}=2$ because the MPS with the bond dimension of $2$ can be exactly implemented on a quantum circuit [44].
To simulate the heat equation (53), we set $n_{0}=n_{1}=4$, $\tau=1.0\times 10^{-1}$, $T=10$, $h=1$, and $\kappa=0.1$. We simulated quantum circuits by the state vector calculation. The number of qubits for encoding this problem is $n_{0}+n_{1}+n_{\mathrm{anc}}=16$. As an initial state, we prepared
@@
This dynamics of the temperature describes how the initial temperature diffuses as time passes, which is a non-unitary dynamics. Note again that such direct visualization requires the full quantum state tomography and thus is impractical for a real quantum computer. Furthermore, we need to repeatedly run quantum circuits to successfully measure all zeros of the ancilla register when implementing the proposed algorithm on the actual quantum computer. In this work, we performed all the computation by statevector simulatior to verify the validity of the proposed method. We would like to conduct our future work that takes into account the sample complexity of required measurements for successfully extracting some meaningful information from the prepared time-evolved states. When solving the heat equation, one might be interested in estimating the maximum temperature in the computational domain, which corresponds to the maximum amplitude of the time-evolved quantum state. Although exactly finding the maximum amplitude is a difficult problem because of the exponentially large Hilbert space, it is easy to get an idea of whether the temperature is concentrated in a specific area or is more uniform throughout the domain by measuring multiple times in the computational basis. Then, we can also estimate the temperature of the identified regions using the amplitude estimation algorithm. Also, the demonstration shown in the current work involved the two systems of PDE: one is a conservative with spatially varying parameters, and the other is non-conservative with uniform parameters, and did not tackle with non-conservative systems of PDEs with spatially varying parameters due to the computational cost although the formulation and algorithm of our proposed method include such systems. We also address this in our future work.
@@ V Conclusion
In the present study, we focused on PDEs without source terms (i.e., $\bm{b}(t)$ in Eq. (1)) for simplicity, while the original LCHS method can deal with equations including the source term, as shown in Eq. (2). Since PDEs with source terms are essential for practical applications in CAE, we plan to address this problem in our future work. Also, we would like to investigate important observables in more detail for a specific problem setting. For example, in addition to the local quantities at a specific region discussed in Eq. (52), global (averaged) quantities such as effective stiffness or conductivity of a composite materials can be promising options in material science. Additionally, our future work will tackle the simulation of non-linear PDEs, which have already been discussed theoretically in the literature [51, 52]. We believe our current work will advance the applicability of quantum computers in the field of CAE.
@@ Acknowledgment
Appendix B Changes in the norm by PDEs with the first-order derivative in time
Since the success probability of solving PDEs by LCU depends on the changes in the norm over the simulation time, we would like to evaluate the changes in the norm by PDEs with the first-order derivative in time as well. The time derivative of the squared norm of the field $u(t,\bm{x})$ is calculated as
$\int_{\Omega}\frac{\partial u(t,\bm{x})^{\dagger}u(t,\bm{x})}{\partial t}\mathrm{d}\bm{x}$
$=\int_{\Omega}u(t,\bm{x})^{\dagger}\left(\nabla\cdot\kappa\nabla-\bm{\beta}\cdot\nabla-\alpha\right)u(t,\bm{x})\mathrm{d}\bm{x}+c.c$
$=\int_{\partial\Omega}u(t,\bm{x})^{\dagger}\bm{n}\cdot\kappa\nabla u(t,\bm{x})\mathrm{d}\bm{x}-\int_{\Omega}\left(\nabla u(t,\bm{x})\right)^{\dagger}\cdot\kappa\nabla u(t,\bm{x})\mathrm{d}\bm{x}$
$\quad-\int_{\Omega}u(t,\bm{x})^{\dagger}\bm{\beta}\cdot\nabla u(t,\bm{x})\mathrm{d}\bm{x}-\int_{\Omega}\alpha u(t,\bm{x})^{\dagger}u(t,\bm{x})\mathrm{d}\bm{x}+c.c.$
$=-2\int_{\Omega}\left(\nabla u(t,\bm{x})\right)^{\dagger}\cdot\kappa\nabla u(t,\bm{x})\mathrm{d}\bm{x}-2\int_{\Omega}\alpha\left|u(t,\bm{x})\right|^{2}\mathrm{d}\bm{x}$
$\quad-\int_{\partial\Omega}\left|u(t,\bm{x})\right|^{2}\bm{n}\cdot\bm{\beta}\mathrm{d}\bm{x}+\int_{\Omega}\left|u(t,\bm{x})\right|^{2}\nabla\cdot\bm{\beta}\mathrm{d}\bm{x},$ (62)
where the first term in the third line integrating over the boundary $\partial\Omega$ vanishes owing to the boundary condition. In the last equality, we applied the integration by parts for the complex conjugate term of the third term, which is about the advection. This equation implies that the norm of the field $u(t,\bm{x})$ monotonically decreases over the time evolution by the PDE (16) when the velocity field $\bm{\beta}(\bm{x})$ satisfies the condition that the last two terms vanish, i.e., the conservative law $\nabla\cdot\bm{\beta}(\bm{x})=0$ and $\bm{n}\cdot\bm{\beta}(\bm{x})=0$. When such a condition is satisfied, the diffusion and absorption coefficients $\kappa(\bm{x})$ and $\alpha(\bm{x})$ determine the decrease of the norm and consequently the success probability of LCU.
Appendix C Acoustic equation